150502 UCSC

Had fan on LML system (RH 91) set backwards – so no sample even though doors opened.  No apparent wetness on ground in morning. UCSC rain sensor system did not open.  Tested both systems and they work properly.  Rinsed out LML collector and jar with DI and reinstalled.

150501 UCSC

Odd post sunrise fog moved in after a really hot day.

Running 2 fog collectors on ISB roof.  One has the rain sensor with mini-collector funnel, connected to UCSC network (email me for the http address to access weather data and images of jar).  The other has the old 2014 Fogon.py program with 91% RH cutoff.

The 91% RH triggered the 2014 system (doors open).

Downloading the passive fog data from the spectrum data loggers.

Begin blank checking intensive